# Rearrangements of distributions on integers that minimize variance

### v1

@@ Abstract
Let $X$ be a random variable on a set of integers $\{x_{1},\dots,x_{N}\}$ such that $\mathbb{P}(X_{i}=x_{i})=p_{i}$, $i\in\{1,\dots,N\}$. Let $(p^{(1)},\dots,p^{(N)})$ be the sequence $(p_{1},\dots,p_{N})$ ordered non-increasingly. Let $X^{+}$ be the random variable defined by $\mathbb{P}(X=0)=p^{(1)}$, $\mathbb{P}(X=1)=p^{(2)}$, $\mathbb{P}(X=-1)=p^{(3)},\dots,\mathbb{P}(X=(-1)^{N}\lfloor\frac{N}{2}\rfloor)=p^{(N)}$. In this short note we generalize and prove the inequality ${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$.
@@ 1 Introduction
The question whether (2) always holds arose while applying a similar argument in [3]. In the present short note we provide a straightforward proof of (2) as we were not able to find it mentioned in the literature.
@@
Let $X$ be a random variable. Let $f:[0,+\infty)\to[0,+\infty)$ be a non-decreasing function such that $\inf_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|)<\infty$. Define a number
$D_{f}(X):=\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
And the set
$M_{f}(X):=\arg\min_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|).$
@@
Let $X$ be a random variable supported on a finite set of integers. Assume that $f:[0,+\infty]\to[0,+\infty)$ is non-decreasing. Then
@@
Furthermore, suppose that $f(x)$ has a positive derivative for $x>0$ and a right derivative at 0 such that $f^{\prime}(0+)=0$. Then (4) is strict unless XkX-k is distributed as $X^{+}$ or $X^{+}$ for some integer $k$.
@@ 2 Proofs
Let $a\in M_{f}(X)$. We denote $D_{f}(X)={\bf p}\cdot{\bf v}$ where ${\bf p}=(p^{(1)},p^{(2)},\ldots,p^{(N)})$ and
@@
This finishes the proof of (4).
Assume now the additional properties of $f$ stated in the second part of the theorem. Assume $D_{f}(X)=D_{f}(X^{+})$, but $X$ is not a translation of $X^{+}$ or $X^{+}$. We will follow the proof of (4) and obtain a contradiction.
@@
Suppose first that $a^{\prime}\not\in\{0,\frac{1}{2}\}$. Then, since $f$ is strictly increasing for $x>0$, identity is the unique permutation that orders $(x_{i})$ non-decreasingly. When $a\in(0,\frac{1}{2})$ this corresponds to placing the probabilities $p^{(1)},\dots,p^{(N)}$ on $0,1,-1,\dots$ respectively as in the distribution of $X^{+}$. Similarly, when $a\in(\frac{1}{2},1)$, this corresponds to placing them on $1,0,2,-1,\dots$ respectively as in the distribution of $1-X^{+}$.

### v2

@@ Abstract
Let $X$ be a random variable on a set of integers $\{x_{1},\dots,x_{N}\}$ such that $\mathbb{P}(X_{i}=x_{i})=p_{i}$, $i\in\{1,\dots,N\}$. Let $(p^{(1)},\dots,p^{(N)})$ be the sequence $(p_{1},\dots,p_{N})$ ordered non-increasingly. Let $X^{+}$ be the random variable defined by $\mathbb{P}(X^{+}=0)=p^{(1)}$, $\mathbb{P}(X^{+}=1)=p^{(2)}$, $\mathbb{P}(X^{+}=-1)=p^{(3)},\dots,\mathbb{P}(X^{+}=(-1)^{N}\lfloor\frac{N}{2}\rfloor)=p^{(N)}$. In this short note we generalize and prove the inequality ${\mathrm{Var}}\,X^{+}\leq{\mathrm{Var}}\,X$.
@@ 1 Introduction
The question whether (2) always holds arose while applying a similar argument in [3]. In the present short note we provide a straightforward proof of (2) as we were not able to find it mentioned in the literature. A special case of the present result is used, along with many other ideas, in the proof of a much more general result of [3].
@@
Let $X$ be a random variable. Let $f:[0,+\infty)\to[0,+\infty)$ be a non-decreasing function. Define a number
$D_{f}(X):=\inf_{a\in\mathbb{R}}{\mathbb{E}}\,f(|X-a|)$
and, in the case $D_{f}(X)<\infty$, the set
$M_{f}(X):=\{a\in\mathbb{R}:{\mathbb{E}}\,f(|X-a|)=D_{f}(X)\}.$
@@
Let $X$ be a random variable supported on a finite set of integers. Assume that $f:[0,+\infty)\to[0,+\infty)$ is non-decreasing and $D_{f}(X)<\infty$. Then
@@
Furthermore, suppose that $f$ is continuous with a positive derivative for $x>0$ and a right derivative at 0 such that $f^{\prime}(0+)=0$. Then (4) is strict unless XkX-k is distributed as $X^{+}$ or $X^{+}$ for some integer $k$.
@@ 2 Proofs
It can be checked that $M_{f}(X)$ is non-empty when $f$ is continuous. First assume that $M_{f}(X)\neq\emptyset$. Let $a\in M_{f}(X)$. We denote $D_{f}(X)={\bf p}\cdot{\bf v}$ where ${\bf p}=(p^{(1)},p^{(2)},\ldots,p^{(N)})$ and
@@
This finishes the proof of (4) when $M_{f}(X)\neq\emptyset$. For the general case, by definition, for any $\epsilon>0$ we can find $a=a(\epsilon)$ such that ${\mathbb{E}}\,f(|X-a|)\leq D_{f}(X)+\epsilon$. The same argument as above shows that $D_{f}(X)\geq D_{f}(X^{+})-\epsilon$. Since $\epsilon>0$ is arbitrary, we conclude that $D_{f}(X)\geq D_{f}(X^{+})$.
Assume now the additional properties of $f$ stated in the second part of the theorem. Now $f$ is continuous, so $M_{f}(X)\neq\emptyset$. Assume $D_{f}(X)=D_{f}(X^{+})$, but $X$ is not a translation of $X^{+}$ or $X^{+}$. We will follow the proof of (4) and obtain a contradiction.
@@
Suppose first that $a^{\prime}\not\in\{0,\frac{1}{2}\}$. Then, since $f$ is strictly increasing for $x>0$, identity is the unique permutation that orders the components of $\bf v$ non-decreasingly. When $a\in(0,\frac{1}{2})$ this corresponds to placing the probabilities $p^{(1)},\dots,p^{(N)}$ on $0,1,-1,\dots$ respectively as in the distribution of $X^{+}$. Similarly, when $a\in(\frac{1}{2},1)$, this corresponds to placing them on $1,0,2,-1,\dots$ respectively as in the distribution of $1-X^{+}$.